OpenStudy (anonymous):

Find the approximate value of the circumference of a circle with the given radius. Use = 3.14. Round your answer to the nearest thousandth. 3.21 ft.

OpenStudy (anonymous):

use the formula for the circumference 2 x pi x radius

OpenStudy (anonymous):

circumference = 2 π r = 2 * 3.14 * 3.21 = 20.159 ft

OpenStudy (anonymous):

If u use π as 22/7, then circumference = 2 π r 22 = 2 * ----- * 3.21 7 = 20.177 ft